# Think2Drive: Efficient Reinforcement Learning by Thinking in Latent World Model for Quasi-Realistic Autonomous Driving (in CARLA-v2)

### v1

@@ 6.2 Simulator Execution Mode
6.3 Terminal Condition.
To save the training time cost, we set several termination conditions to terminate episodes in advance. An episode terminates if and only if one of these conditions is triggered:
1) Route completed. The ego vehicle has finished all its target routes.
2) Route deviation. The ego vehicle deviates from its target route far beyond 8m.
3) Collision. The ego vehicle collides with vehicles, walkers, or other static obstacles.
4) Timeout. The ego hasn’t moved for over 50 seconds.
5) Traffic light/sign infraction. The ego vehicle runs a stop sign or a red light.
6.4 Reward Shaping.
Our reward is shaped to make the planner keep safe driving and finish the target route as much as possible, which consists of 4 parts:
1) Speed reward $r_{speed}$ is used to train the ego vehicle to keep a safe speed, depending on the distance to other objects and their type.
2) Travel reward $r_{travel}$ is the distance traveled along the target routes at each tick of CARLA. Travel reward encourages the ego vehicle to finish more target routes.
3) Deviation penalty $p_{deviation}$ is the negative value of the distance between the ego vehicle and the lane center. It is normalized by the max deviation threshold $D_{max}$.
4) Steering cost $c_{steer}$ is used to make the ego vehicle drive more smoother. We set it as a constant negative value and send it to the ego whenever the current steering is different from the last one.
The overall reward function is defined as follows:
$r=$ $\alpha_{sp}r_{speed}+\alpha_{tr}r_{travel}$ (7)
$+\alpha_{de}p_{deviation}+\alpha_{st}c_{steer}$
6.5 Model Size
Table 7: Hyper-parameters of the neural network. The CNN encoder maps the $\times 128$ 128 BEV to the $\times 4$ 4 feature map by $\times 4$ 4 convolutional kernel with stride 2. The flattened feature maps, concatenated with the state features output by the MLP encoder, are then input to the RSSM which consists of GRU cells and a few dense layers. The structure of the decoder inverts that of the encoder and outputs the reconstructed mask and state vector. Please refer to DreamerV3 [17] for details about network.
GRU
units
CNN
multiplier
Dense
hidden units
MLP
layers
Params
@@ 6.5 Model Size
6.6 Hyperparameters of Training
Table 8: Hyperparameter settings of training
Hyperparameters Notation Value
      General
Replay capacity (FIFO) 3e5
Batch size $B$ 16
Batch length $T$ 64
Activation LayerNorm + SiLU
Reset step 0.8M
    World Model Learning
World model train ratio 16
Reward prediction loss scale 10.0
Number of latents 32
Classes per latent 32
Reconstruction loss scale $\beta_{pred}$ 1.0
Dynamics loss scale $\beta_{dyn}$ 0.5
Representation loss scale $\beta_{rep}$ 0.1
Learning rate 1e-4
Adam epsilon $\epsilon$ 1e-8
Gradient clipping 1000
    Planner Learning
Planner train ratio 16, 32, 128, 256
Imagination horizon 15
Discount horizon $1/(1-\gamma)$ 333
Return lambda $\lambda$ 0.95
Critic EMA decay 0.98
Critic EMA regularizer 1.0
Return normalization scale $S$ $Per(R,95)-Per(R,5)$
Return normalization limit 1
Return normalization decay 0.99
Actor entropy scale 3e-4
Learning rate 3e-5
Adam epsilon $\epsilon$ 1e-8
Gradient clipping 100
Speed reward scale $\alpha_{sp}$ 1.0
Travel reward scale $\alpha_{tr}$ 1.0
Deviation reward scale $\alpha_{de}$ 2.0
Steer cost scale $\alpha_{st}$ 0.5

### v2

@@ Appendix 0.C Simulator Execution Mode
Appendix 0.D Hyper-parameters
Table 7: Hyper-parameters of the neural network. The CNN encoder maps the $128\times 128$ BEV to the $4\times 4$ feature map by $4\times 4$ convolutional kernel with stride 2. The flattened feature maps, concatenated with the state features output by the MLP encoder, are then input to the RSSM which consists of GRU cells and a few dense layers. The structure of the decoder inverts that of the encoder and outputs the reconstructed mask and state vector. Please refer to DreamerV3 [16] for details about network.
GRU recurrent units CNN multiplier Dense hidden units MLP layers Parameters